# DDOS: The Drone Depth and Obstacle Segmentation Dataset

### v1

@@ 5 Dataset Statistics
In this section, we provide a comprehensive analysis of key properties inherent in the DDOS dataset. The visual representation in Figure 2 vividly illustrates the distribution of annotations across diverse classes within DDOS. Significantly, the dataset adeptly captures and represents various classes of thin objects, even when these objects occupy a relatively small number of pixels in each image. This nuanced representation ensures that DDOS offers a substantial and well-balanced dataset for thin object classes. This richness in diversity is paramount for facilitating thorough analysis, robust algorithm training, and effective evaluation, particularly in addressing the challenges associated with thin objects in real-world scenarios. The carefully crafted distribution of classes within DDOS contributes to its utility as a reliable benchmark for advancing the capabilities of algorithms designed for thin object detection and segmentation.
Continuing our exploration, we delve into the examination of pitch and roll angles recorded during the flights. As depicted in Figure 3, the visualization illustrates a broad spectrum of pitch and roll angles, highlighting substantial variations in the drone’s orientation throughout the dataset. Remarkably, despite the drone’s predominant forward movement in the dataset, the recorded angles exhibit noteworthy diversity. This diversity in orientation contributes invaluable viewing perspectives, offering a holistic evaluation of algorithms under diverse flight conditions. The extensive distribution of pitch and roll angles underscores the DDOS dataset’s capacity to emulate real-world scenarios, where drones navigate through varying pitch and roll angles. This feature ensures the dataset’s efficacy in training and assessing algorithms to perform reliably across a spectrum of orientation challenges encountered during drone flights.
@@
To gain an intuitive understanding of the spatial distribution of flight paths within a singular environment, we visually present a subset of the recorded trajectories in Figure 4. The depicted flight paths showcase a diverse array of patterns, ranging from sharp turns and straight lines to curved trajectories. These variations authentically capture the complexity and dynamic nature of the simulated environments. Furthermore, an overhead view of the relative flight paths, presented in Figure 5, offers a normalized perspective with a common starting point and direction. This visualization emphasizes the diverse flight trajectories and patterns observed across individual flights, providing a comprehensive overview of the spatial dynamics inherent in the DDOS dataset. Such a representation is instrumental in offering insights into the intricate navigation challenges that algorithms must address, reinforcing the dataset’s efficacy in training and evaluating models under diverse and realistic conditions.
@@
Expanding our analysis, we delve into the distributions of altitude and speed during the flights, along with the distribution of depth recorded in the depth maps, as illustrated collectively in Figure 6. Examining the altitude distribution reveals that the drone operates at varying heights, encompassing low-level flights near the ground to higher altitudes. The distribution of speed elucidates a spectrum of velocities encountered during the flights, showcasing diverse flight behaviors and maneuvering speeds. Moreover, the depth distribution offers insights into the range and distribution of depth values recorded in the depth maps, shedding light on the variations in perceived depth across the dataset. This multifaceted exploration of altitude, speed, and depth distributions provides a comprehensive understanding of the dynamic spatial characteristics inherent in the DDOS dataset, crucial for informing the development and evaluation of algorithms designed to navigate and perceive diverse environments encountered by drones.
@@ 6 Depth metrics
In this formula, represents the absolute relative depth estimation error for the ultra thin class. denotes the total number of samples in the ultra thin class. and represent the ground truth depth and estimated depth for the -th pixel sample, respectively. The formula calculates the average absolute relative difference between the ground truth and estimated depths for all samples in the ultra thin class.
By evaluating the absolute relative depth error for this class, we can assess the performance of depth estimation algorithms in accurately capturing the depth information of ultra thin objects, such as wires and cables. The introduction of class-specific depth metrics enables a more comprehensive evaluation of depth estimation algorithms tailored to the unique challenges posed by different object categories encountered in drone views.
Trivially, extending this approach to all classes, the general formula for class-specific depth metrics becomes:
By assessing the absolute relative depth error for each class, we gain insights into the performance of depth estimation algorithms in accurately capturing the depth information of specific objects, such as wires and cables. The incorporation of class-specific depth metrics facilitates a more nuanced evaluation of depth estimation algorithms, addressing the unique challenges presented by various object categories encountered in drone views.
@@
BinsFormer [8] | 0.807 | 0.930 | 0.970 | 0.153 | 7.354 | 0.061 | 0.212 | 19.529 | 1.848 |
SimIPU [7] | 0.811 | 0.931 | 0.969 | 0.149 | 6.366 | 0.056 | 0.206 | 19.470 | 1.770 |
DepthFormer [9] | 0.873 | 0.957 | 0.981 | 0.108 | 5.761 | 0.045 | 0.171 | 16.542 | 1.119 |
@@
BinsFormer [8] | 0.686 | 0.248 | 0.149 | 0.206 | 0.251 | 0.156 | 0.153 | 0.139 | 0.159 | 0.108 |
SimIPU [7] | 0.664 | 0.280 | 0.157 | 0.232 | 0.279 | 0.188 | 0.165 | 0.149 | 0.181 | 0.097 |
DepthFormer [9] | 0.634 | 0.228 | 0.118 | 0.187 | 0.189 | 0.123 | 0.138 | 0.115 | 0.154 | 0.067 |
@@ 7 Baselines
We employ a comprehensive set of metrics to rigorously evaluate the effectiveness of our baselines. These metrics include fundamental measures such as accuracy under the threshold (, ), which assesses the model’s performance within certain proximity thresholds. Additionally, we consider common metrics like mean absolute relative error (AbsRel), mean squared relative error (SqRel), root mean squared error (RMSE), root mean squared log error (RMSElog), and mean log10 error (log10). These metrics provide insights into the accuracy and precision of our depth estimation. To further enhance our evaluation, we integrate the scale-invariant logarithmic error (SILog).
Moreover, in pursuit of a more nuanced evaluation, we leverage our newly proposed suite of metrics known as mean absolute relative class error metrics (). This suite is tailored to assess the performance of our methods at a finer class level, offering a more detailed understanding of their capabilities. This diverse array of metrics ensures a thorough assessment of the selected baselines.
We utilise three different baselines, BinsFormer [8], SimIPU [7] and DepthFormer [9]. BinsFormer proposes a novel framework for monocular depth estimation by formulating it as a classification-regression task, employing a transformer [14] decoder to generate adaptive bins [1]. SimIPU introduces a pre-training strategy for spatial-aware visual representation, utilizing point clouds for improved spatial information in contrastive learning. DepthFormer addresses supervised monocular depth estimation by leveraging a transformer for global context modeling, incorporating an additional convolution branch, and introducing a hierarchical aggregation module. All three baselines demonstrate superior performance in monocular depth estimation compared to state-of-the-art methods across various datasets.
The evaluation of depth estimation baselines is shown in table 2. DepthFormer emerges as the standout performer, consistently surpassing its counterparts in accuracy under different thresholds, mean absolute relative error, and other key metrics. The success of DepthFormer can be attributed to its innovative approach, leveraging a transformer for global context modeling, incorporating additional convolutional branches, and introducing hierarchical aggregation modules. These features enhance both local and global information processing, resulting in superior depth estimation capabilities.
The scrutiny of our novel class-specific depth metrics, as delineated in table 3, unveils the challenges inherent in accurately estimating the depth of specific object classes across all methods under consideration. Notably, the ultra-thin class emerges as a particularly formidable challenge, as evidenced by the discernible difficulties encountered by all methods in providing precise depth estimations for this category. The observed limitations in accurately capturing the depth of ultra-thin objects underscore the complexities associated with fine-grained depth estimation, warranting further investigation and refinement of methodologies. This underscores the necessity for methodological advancements that specifically address drone setting and the intricacies of ultra-thin object depth estimation, signaling a potential avenue for future research and algorithmic refinement. To further highlight these challenges, we present qualitative examples in Appendix C for illustrative purposes.
@@ 8 Conclusion
Accurate depth and semantic segmentation play pivotal roles in diverse computer vision applications, from autonomous navigation to surveillance. However, the shortage of annotated real-world aerial datasets poses a formidable challenge, hindering the development and evaluation of robust models. To address these challenges, this work introduces DDOS, a synthetic dataset tailored for depth and semantic segmentation tasks in aerial views. Leveraging AirSim [12], DDOS serves as a crucial resource for model training, particularly in scenarios where real-world annotated data is scarce. Furthermore, the introduction of novel drone-specific metrics for depth accuracy is a noteworthy contribution, promising to advance the evaluation standards for models operating in drone-specific scenarios.
Inaccuracies in detecting elements such as cables and wires pose significant risks, warranting specialized attention in model development. DDOS and drone-specific metrics introduced in this work not only address current limitations but also pave the way for future advancements in computer vision for aerial views. The tailored metrics offer a nuanced evaluation framework, ensuring that models are proficient not only in general scenarios but also in the intricacies of drone-specific contexts. As such, this research contributes to the broader landscape of safety-conscious drone technologies and fosters advancements in computer vision research for aerial applications.
@@ A.1 Motivation
The Drone Depth and Obstacle Segmentation (DDOS) dataset, was created to address the limitations posed by the scarcity of annotated aerial datasets, specifically for training and evaluating models in depth and semantic segmentation tasks. The primary objective is to focus on the detection and segmentation of thin objects like wires, cables, and fences in aerial views, which are critical for ensuring the safe operation of drones. The dataset aims to fill the gap in existing datasets that predominantly concentrate on common objects and lack representation of fine spatial characteristics of thin objects.
Who created the dataset and on behalf of which entity?
@@ A.2 Composition
Each flight consists of sequential observations, comprising of a high-resolution image captured by a monocular camera affixed to the front of the drone, corresponding depth maps, pixel-level object segmentation masks, optical flow information and surface normals. As well as coordinates, pose and speed information and environment information including weather. All image modalities maintain a resolution of , and the depth maps cover a range from 0 to 100 meters.
@@ A.6 Distribution
Yes, DDOS is hosted on Hugging Face and is openly available:
@@ A.7 Maintenance
RGB Image | Depth Map | Segmentation |
RGB Image | Depth Map | Segmentation |
RGB Image | Ground Truth | BinsFormer [8] | SimIPU [7] | DepthFormer [9] |
@@ Appendix B Additional Examples
We present additional frames from DDOS in Figures 7 and 8. To accentuate finer details, the visualizations focus solely on the RGB image alongside its corresponding depth map and semantic segmentation. The showcased frames exhibit diverse variations in perspective, encompassing different types of vegetation, architectural structures, and meteorological phenomena such as snow cover. Figure 8 highlights examples with instances belonging to the Ultra thin class, as well as examples from the Thin structures and Small mesh classes.
DDOS, serves as a comprehensive resource for the research community, particularly in the domains of depth estimation and segmentation. Its utility is especially evident in scenarios involving aerial perspectives, as encountered by drones, offering valuable insights for discerning thin objects within the visual field.
Appendix C Qualitative Results
In this section, we present qualitative results obtained from the baseline models, as illustrated in Figure 9. A visual inspection reveals substantial similarities among the baseline models in their overall performance. Notably, each struggles when predicting the maximum depth, indicating a shared limitation across the evaluated approaches.
However, a more detailed analysis unveils distinct challenges faced by the BinsFormer model, particularly in its handling of the sky and clouds. This model exhibits a pronounced difficulty, often misjudging the proximity of clouds, thereby leading to inaccurate predictions. On the other hand, all models demonstrate commendable proficiency in accurately predicting thin structures, exemplified by floodlights and electrical poles. Regrettably, a notable deficiency emerges collectively across all models concerning Ultra thin objects, specifically wires, where each model experiences a substantial failure.
This observed limitation underscores the exigency for DDOS. It becomes evident that current state-of-the-art techniques are ill-equipped to effectively address the intricacies associated with ultra-thin objects, emphasizing the imperative for novel methodologies. The introduction of the DDOS dataset is motivated by the need to provide a dedicated resource for advancing research in ultra-thin object depth estimation and segmentation, acknowledging the scarcity of suitable data for addressing this specific challenge.

### v2

@@ 5 Dataset Statistics
In this section, we provide a comprehensive analysis of key properties inherent in the DDOS dataset. Figure 2 illustrates the distribution of annotations across diverse classes within DDOS. Significantly, the dataset adeptly captures and represents various classes of thin structures, even when these objects occupy a relatively small number of pixels in each image. This nuanced representation ensures that DDOS offers a substantial and well-balanced dataset for thin object classes. This richness in diversity is paramount for facilitating thorough analysis, robust algorithm training, and effective evaluation, particularly in addressing the challenges associated with thin structures in real-world scenarios. The carefully crafted distribution of classes within DDOS contributes to its utility as a reliable benchmark for advancing the capabilities of algorithms designed for thin structure detection and segmentation.
In our continued investigation, we analyze the pitch and roll angles observed during flight sessions. As depicted in Figure 3, there is a wide range of pitch and roll angles, indicating significant variations in the drone’s orientation across the dataset. Despite the drone’s primary forward motion, the angles demonstrate a notable diversity. This variety in orientation provides valuable perspectives for evaluating algorithms under different flight conditions. The broad distribution of pitch and roll angles emphasizes the DDOS dataset’s ability to mimic real-world flying scenarios, where drones encounter various orientations. This characteristic enhances the dataset’s utility for training and evaluating algorithms to ensure consistent performance amidst the orientation challenges that drones face in actual flights.
@@
To gain an intuitive understanding of the spatial distribution of flight paths within an environment, we visually present a subset of the recorded trajectories in Figure 4. The depicted flight paths showcase a diverse array of patterns, ranging from sharp turns and straight lines to curved trajectories. These variations authentically capture the complexity and dynamic nature of the simulated environments. Furthermore, an overhead view of the relative flight paths, presented in Figure 5, offers a normalized perspective with a common starting point and direction. This visualization emphasizes the diverse flight trajectories and patterns observed across individual flights, providing a comprehensive overview of the spatial dynamics inherent in DDOS. Such a representation is instrumental in offering insights into the intricate navigation challenges that algorithms must address, reinforcing the dataset’s efficacy in training and evaluating models under diverse and realistic conditions.
@@
Expanding our analysis, we explore the distributions of altitude and speed during the flights, along with the distribution of depth recorded in the depth maps, as illustrated collectively in Figure 6. Examining the altitude distribution reveals that the drone operates at varying heights, encompassing low-level flights near the ground to higher altitudes. The distribution of speed elucidates a spectrum of velocities encountered during the flights, showcasing diverse flight behaviors and maneuvering speeds. Moreover, the depth distribution offers insights into the range and distribution of depth values recorded in the depth maps, shedding light on the variations in perceived depth across the dataset.
@@ 6 Depth Metrics
Here, represents the absolute relative depth estimation error for the Ultra Thin class. denotes the total number of samples (pixels) in the Ultra Thin class, while and represent the ground truth depth and estimated depth for the -th pixel sample, respectively. The formula calculates the average absolute relative difference between the ground truth and estimated depths for all samples in the Ultra Thin class. Trivially, extending this approach to all classes, the general formula for class-specific depth metrics becomes:
Assessing class-specific absolute relative depth errors reveals how well depth estimation algorithms perform, especially for intricate structures like wires and cables. This method offers a detailed evaluation, highlighting how algorithms manage the challenges unique to various structures seen from drone viewpoints. The motivation for this nuanced approach stems from the recognition that traditional metrics fail to adequately represent difficult-to-detect obstacles, such as wires, due to their low pixel count. A thorough investigation into these aspects is essential to accurately gauge the efficacy and robustness of vision systems.
@@
BinsFormer [19] | 0.632 | 0.792 | 0.845 | 0.265 | 16.211 | 0.139 | 0.466 | 38.009 | 6.387 |
SimIPU [18] | 0.760 | 0.918 | 0.964 | 0.225 | 7.095 | 0.070 | 0.245 | 22.715 | 3.302 |
DepthFormer [20] | 0.860 | 0.958 | 0.981 | 0.136 | 5.831 | 0.050 | 0.190 | 18.101 | 1.614 |
@@
BinsFormer [19] | 0.945 | 0.216 | 0.129 | 0.209 | 0.248 | 0.137 | 0.141 | 0.150 | 0.141 | 0.257 |
SimIPU [18] | 1.036 | 0.317 | 0.178 | 0.233 | 0.380 | 0.198 | 0.204 | 0.176 | 0.184 | 0.122 |
DepthFormer [20] | 0.998 | 0.229 | 0.115 | 0.177 | 0.206 | 0.121 | 0.120 | 0.121 | 0.128 | 0.082 |
@@ 7 Baselines
We use a set of commonly-used depth metrics to evaluate the effectiveness of the baselines. These metrics include fundamental measures such as accuracy under the threshold (, ), which assesses the model’s performance within proximity thresholds. Additionally, we use mean absolute relative error (AbsRel), mean squared relative error (SqRel), root mean squared error (RMSE), root mean squared log error (RMSElog), mean log10 error (log10) and scale-invariant logarithmic error (SILog).
Moreover, in pursuit of a more nuanced evaluation, we leverage our newly proposed suite of metrics known as mean absolute relative class error metrics (). This suite is tailored to assess the performance of our methods at a finer class level, offering a more detailed understanding of their capabilities.
We utilize three different baselines, BinsFormer [19], SimIPU [18] and DepthFormer [20]. BinsFormer proposes a novel framework for monocular depth estimation by formulating it as a classification-regression task, employing a transformer [37] decoder to generate adaptive bins [5]. SimIPU introduces a pre-training strategy for spatial-aware visual representation, utilizing point clouds for improved spatial information in contrastive learning. DepthFormer addresses supervised monocular depth estimation by leveraging a transformer for global context modeling, incorporating an additional convolution branch, and introducing a hierarchical aggregation module.
When evaluated using standard depth metrics, the baselines exhibit satisfactory performance, as shown in Table 2. However, using our class-specific depth metrics, shown in Table 3 and depicted in Figure 7, unveils substantial challenges in achieving accurate depth estimations for certain object classes. Specifically, the Ultra Thin category is exceptionally challenging, with all tested methods failing to provide accurate depth estimations.
These findings highlight the importance of developing methodologies that are specifically tailored to enhance depth estimation accuracy for ultra-thin structures, particularly in drone-based applications. Future research should focus on addressing these challenges, aiming to enhance the precision and reliability of depth estimations for these challenging scenarios.
@@ 8 Conclusion
In summary, we introduce the DDOS dataset along with novel drone-specific depth metrics, marking a pivotal advancement in the field of autonomous drone navigation. The DDOS dataset addresses the critical challenges of detecting thin structures and operating under varied weather conditions, thereby filling an essential gap in the current scope of drone research. Through a detailed analysis of the dataset and the deployment of tailored evaluation metrics, we provide a nuanced methodology for systematically assessing the performance of depth estimation algorithms in drone-specific scenarios.
These efforts establish a new standard for future investigations aimed at enhancing the safety and efficiency of drone navigation through superior depth estimation and semantic segmentation techniques. The introduction of the DDOS dataset and corresponding metrics not only propels forward the development of drone technology but also extends the potential for computer vision applications within aerial environments. Our work lays a crucial groundwork for future innovations, steering the creation of algorithms that adeptly navigate the complexities of real-world settings, thus amplifying the functional prowess of drones across a multitude of industries.
@@ A.1 Motivation
The Drone Depth and Obstacle Segmentation (DDOS) dataset, was created to address the limitations posed by the scarcity of annotated aerial datasets, specifically for training and evaluating models in depth and semantic segmentation tasks. The primary objective is to focus on the detection and segmentation of thin structures like wires, cables, and fences in aerial views, which are critical for ensuring the safe operation of drones. The dataset aims to fill the gap in existing datasets that predominantly concentrate on common structures and lack representation of fine spatial characteristics of thin structures.
Who created the dataset?
@@ A.2 Composition
Each flight consists of sequential observations, comprising of a high-resolution image captured by a monocular camera affixed to the front of the drone, corresponding depth maps, pixel-level object segmentation masks, optical flow information and surface normals. As well as coordinates, pose and speed information and environment information including weather. All image modalities maintain a resolution of , and the depth maps cover a range from .
@@ A.6 Distribution
Yes, DDOS is hosted on Hugging Face and is available at:
@@ A.7 Maintenance
Image Depth Segmentation
@@ Appendix B Additional Examples
In this section, we present further examples from the DDOS dataset, as illustrated in Figures 9 and 8. These examples are specifically selected to highlight the dataset’s diversity and the intricate details captured within. For clarity and emphasis on these finer aspects, the visualizations are confined to the RGB images, accompanied by their respective depth maps and semantic segmentations. Notably, Figure 9 offers a glimpse into the diverse perspectives encompassed within DDOS. Conversely, Figure 8 is dedicated to showcasing scenarios captured during low altitude flights in snowy conditions, underscoring the dataset’s versatility and the challenging environments it encompasses.
DDOS, serves as a comprehensive aerial resource for the research community, particularly in the domains of depth estimation and segmentation. Its utility is especially evident in scenarios involving aerial perspectives, as encountered by drones, offering valuable insights for discerning thin structures within the visual field.